An object that travels in a straight line and is operated by remote control moves away from the starting point by a distance of 125m in 12.5s and returns to half the distance traveled from the starting point 14.2s. Calculate the average velocity.
Answer #1

Average velocity V ,= total distance/total time

=) V = (125+62.5)/(12.5+14.2) = 187.5/26.7 = 7.03m/s

Hence average velocity is 7.03m/s